Suppose the c.d.f. of X is F(t) 3 for 0<t< (a) What is F(5)? (b) What is F(-5)? (c) Compute the p.d.f of X. (d) Compute the mean of X (e) Compute the variance of X. (f) Compute the standard deviation of X (g) Compute the squared coefficient of variation of X.
Answer #1

a)F(5)=1 as 5 is above upper bound 1.

b) F(-5)=0 as -5 is below lower bound of 0

c)

pdf fo X : f(x) =(d/dx)F(x)=(d/dx)*t3 =3t2

d)

mean of X =E(X)=\int_{0}^{1} t*f(t) dt =\int_{0}^{1}3t3 dt =3t4/4 |10 =3/4

e)

E(X2)=\int_{0}^{1} t2*f(t) dt =\int_{0}^{1}3t4 dt =3t5/5 |10 =3/5

Variance of X =Var(X)=E(X2)-(E(X))2 =3/5-(3/4)2 =3/80

f)

std deviaiton of X =sqrt(Var(x))=sqrt(3/80)

g)squared coefficent of variation =Var(x)/(E(X))2 =(3/80)/(9/16)=1/15 or 6.67%